OpenStudy (vibeyyy.waves):

Line L is perpendicular to the line with equation x + 2y = 8 and passes through the origin. Determine the equation of line L.

OpenStudy (mathmale):

Line L is perpendicular to line x + 2y = 8. 1. Find the slope of x + 2y = 8. 2. Find the negative reciprocal of that slope. The result is the slope of your new line. 3. Your new line passes through the origin. This implies that it passes through (0,0). 4. Write out the equation y=mx + b. Substitute your new slope for m here. Subst. 0 for x and 0 for y. Find b. 5. Re-write your equation in the form Ax + By = C. Your job, at this point, is basically to determine the value of the constant, C.